OpenStudy (anonymous):

PLEASE HELP ME !! What is the equation in standard form of the line which passes through (4, −2) and has a slope of −3?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@jdoe0001

OpenStudy (phi):

Notice they give you a point and a slope. Do you know the point-slope formula for a line?

OpenStudy (anonymous):

x2-x1 y2-y1 ?

OpenStudy (phi):

that looks like how to find the slope (but it's upside down) see http://www.freemathhelp.com/point-slope.html

OpenStudy (phi):

once you can write down the equation in point slope form, you then change it into standard form the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

OpenStudy (anonymous):

how do you put it in point slope form?

OpenStudy (phi):

did you look at the site posted up above?

OpenStudy (anonymous):

i can't open it

OpenStudy (phi):

\[ y - y_1 = m(x-x_1) \]

OpenStudy (anonymous):

-2 -2 = -3(4-4) ?

OpenStudy (phi):

the plain x and y are left as letters. the \( (x_1,y_1) \) stand for the x and y *numbers* that they give you

OpenStudy (anonymous):

y - (-2) = -3(x - 4)

OpenStudy (phi):

yes. The next step is change it into Ax + By = C where A is a positive integer, and B, and C are integers. the first step is distribute the -3 (multiply all the terms inside the parens by -3) also change y - -2 into y+2

OpenStudy (anonymous):

y + 2 = -3x +12

OpenStudy (phi):

now add +3x to both sides. (put the x term in front of the y term because we want the answer to match with Ax + By = C )

OpenStudy (phi):

then add -2 to both sides.

OpenStudy (anonymous):

3x + y = 10

OpenStudy (anonymous):

thats correct?

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you so much

OpenStudy (phi):

yw